# Exploiting Task Relationships in Continual Learning via Transferability-Aware Task Embeddings

### v3

@@ Abstract
Continual learning (CL) has been a critical topic in contemporary deep neural network applications, where higher levels of both forward and backward transfer are desirable for an effective CL performance. Existing CL strategies primarily focus on task models — either by regularizing model updates or by separating task-specific and shared components — while often overlooking the potential of leveraging inter-task relationships to enhance transfer. To address this gap, we propose a transferability-aware task embedding, termed H-embedding, and construct a hypernet framework under its guidance to learn task-conditioned model weights for CL tasks. Specifically, H-embedding is derived from an information theoretic measure of transferability and is designed to be online and easy to compute. Our method is also characterized by notable practicality, requiring only the storage of a low-dimensional task embedding per task and supporting efficient end-to-end training. Extensive evaluations on benchmarks including CIFAR-100, ImageNet-R, and DomainNet show that our framework performs prominently compared to various baseline and SOTA approaches, demonstrating strong potential in capturing and utilizing intrinsic task relationships. Our code is publicly available at https://github.com/viki760/H-embedding-Guided-Hypernet.
@@ 1 Introduction
Continual learning (CL), also referred to as incremental learning or life-long learning, has become an essential topic in the modern application of deep neural networks. In CL settings, a model is expected to sequentially learn a series of tasks (typically involving either category shifts or data distribution changes (qu2021recent)) to progressively enhance its capabilities (wang2024comprehensive). A key practical challenge in this process is catastrophic forgetting (CF) (kirkpatrick2017overcoming), a phenomenon where learning a new task undermines the knowledge acquired from previous ones.
@@
The mitigation of CF has long been a primary focus in CL research (wang2024comprehensive) due to its severe hindrance to the overall accumulation of model capacity. Existing approaches can be broadly categorized into three classes: rehearsal-based methods (e.g., ER (robins1995catastrophic), DER (buzzega2020dark)) maintain a memory buffer of previous samples for later replay; regularization-based methods (e.g., LwF (li2017learning), SI (zenke2017continual)) constrain parameter updates to protect prior knowledge; and architecture-based methods (e.g., PackNet (mallya2018packnet), WSN (kang2022forget)) allocate task-specific and task-sharing components of the model from different architecture levels.
@@
However, strategies that explicitly alleviate forgetting often compromise the acquisition of new knowledge, prioritizing retention over adaptability. In contrast, an effective CL system should facilitate the intrinsic knowledge transfer across all tasks, improving future task performance via forward transfer, and preserving (or even enhancing) prior task performance via backward transfer (von2020continual). Moreover, most existing works adopt a model-centric perspective [^1] More discussion in Sec. 2., relying on posterior information obtained during training to manage cross-task interactions. Yet in CL settings, task sequences often exhibit certain degrees of noise and diversity, making such approaches sensitive to sample- and task-level variations. This can lead to potential negative transfer, reduced robustness, and limited scalability, particularly as the number or complexity of tasks increases.
@@
These observations point to a fundamental question in continual learning: Robust and positive knowledge transfers across tasks rely on capturing the underlying task relationships, but how do we efficiently learn and utilize such relations in CL settings? Despite the valuable contributions of previous explorations (e.g. jin2022helpful; riemer2018learning), they mostly remain constrained by a posterior model-centric paradigm with its associated drawbacks, leaving the prior exploitation of task relations largely underexplored. Based on this insight, we argue that the incorporation of a prior, task-relation-aware guidance can effectively leverage task relation information and help mitigate instability and performance degradation in CL, especially over long or challenging task sequences.
@@
Therefore, in this work we explicitly bring focus to the prior exploitation of task relationships by introducing a statistically grounded, task-relation-aware embedding, and propose a CL framework under its guidance. Specifically, recognizing the shared goal between identifying inter-task relationships and the role of transferability metrics (ding2024model) in evaluating source-target task compatibility, we propose an online embedding scheme named H-embedding, which distills task-level transferability into a low-dimensional representation through optimization before task training. H-embedding can be learned efficiently without revisiting previous samples by maximizing the consistency between the Euclidean distances among embeddings and the H-score transferability (bao2019information) among corresponding tasks. To ensure alignment between embedding distances and transferability scores, we apply analytic hierarchy process (AHP) normalization, which significantly improves learning stability over long task sequences while maintaining efficiency. Built upon this embedding, we present a hypernetwork-based CL framework, where a task-conditioned hypernetwork generates task-specific model weights based on H-embedding modulated task embeddings.
@@ 2.1 Continual Learning Strategies
Up to now, there have been many studies dedicated to CL strategies, most of them involving the rehearsal of previous data to alleviate the knowledge degradation caused by CF (e.g. robins1995catastrophic; buzzega2020dark). However, growing privacy and data safety concerns have made this solution not always feasible, bringing increased attention to the rehearsal-free CL setting smith2023closer. Most rehearsal-free methods rely on regularization—either in parameter space (e.g. kirkpatrick2017overcoming; zenke2017continual) or feature space (e.g. li2017learning; rebuffi2017icarl)—by constraining model updates to preserve past knowledge. While effective in CF mitigation, these methods are often based on assuming task similarity and pose limitations to model adaptability. Architecture-based approaches (wang2024comprehensive) offer an alternative by separating task-specific and shared components at various architectural levels (e.g. mallya2018packnet; wortsman2020supermasks; jin2022helpful). With the rise of pretrained models, parameter efficient finetuning (PEFT) components like prompts (e.g. wang2022learning; smith2023coda; wang2023hierarchical) and LoRA modules (e.g. liang2024inflora; wu2025sd) are notably gaining popularity in this paradigm. Nevertheless, the allocation of model parts to different tasks usually comes with scaling problems with the growth of task numbers, potentially resulting in insufficient model capacity or excessive model size growth. Task transferability has recently been explored to enhance CL performance (ermis2022memory). Yet, the current method relies on storing both past models and selected samples, which is incompatible with the rehearsal-free setting. Developing strategies that effectively incorporate task transferability while remaining compatible with the online CL paradigm therefore remains an open question.
@@ 2.2 Transferability Metrics
Task transferability (zamir2018taskonomy) investigates the relationships between tasks and provides an effective method to evaluate and select source tasks in transfer learning. It also plays a crucial role in developing strategies for multi-task learning and meta-learning. For ease of use, previous studies have proposed metrics based on task models and data distributions for a quick estimation of transferability (ding2024model). H-score (bao2019information; ibrahim2022newer; wu2024h) uses an information-theoretic framework to evaluate transferability by solving a maximum correlation problem. NCE (tran2019transferability) employs conditional entropy to assess transferability and task difficulty. LEEP score (nguyen2020leep; agostinelli2022transferability) offers a more generalized metric, defined by measuring the performance of a classifier developed from source model predictions when applied to the target task. LogME (you2021logme) assesses target task accuracy using a formulation integrating all possible linear classifiers derived from source model features. OTCE (tan2021otce; tan2024transferability) combines optimal transport with conditional entropy to both estimate the domain and task difference between source and target. These metrics are mostly designed with differed assumptions and source accessibility, with their use applicable to different problem settings.
@@ 3.2 Hypernets
Hypernets (ha2017hypernetworks), or hypernetworks, are specialized neural networks that generate weights of another neural network, i.e., the target network. Given a target model or component $f$, a hypernetwork $f_{h}$ produces its parameters $\Theta$ conditioned on an auxiliary input. This input can be naturally interpreted as a task embedding, as it reflects the task-specific information needed to generate appropriate parameters. Denoting the task embedding as $e$ and hypernet parameters as $\Theta_{h}$, we have $\Theta=f_{h}(e,\Theta_{h})$. By leveraging task embeddings, hypernets enable flexible parameter modulation across tasks, offering a mechanism for task adaptation without direct weight sharing or parameter interference. Recently, hypernets have gained recognition as a potent tool in deep learning, proving effective across diverse deep learning tasks (chauhan2023brief), including continual learning (von2020continual), causal inference (chauhan2024dynamic), domain adaptation (volk2022example), few-shot learning (sendera2023hypershot), and reinforcement learning (sarafian2021recomposing).
@@ 3.3 H-score
H-score is firstly introduced by huang2019information in huang2019information as a metric assessing the informativeness of features for a task. Theoretically derived from the maximal correlation interpretation of deep neural networks, its mathematical foundation roots to the information theory work known as maximal correlation analysis, which originates from the works of Hirschfeld, Gebelein and Renyi (hirschfeld1935connection; gebelein1941statistische; renyi1959measures) and has been followed and further explored by a broad spectrum of successive work. The H-score of $f$ with regard to the task casting $X$ to $Y$ is defined as:
@@
with input data $X$, label $Y$ and feature extractor $f(X)$. Subsequent research has extended the application of H-score, establishing it as a metric for transferability and validating its efficiency through extensive experimental evaluations (bao2019information; ibrahim2022newer). This underscores H-score’s potential for transfer learning and its relevance to related challenges. We have chosen to incorporate H-score into our framework due to its robust theoretical reliability, alignment of assumptions with our problem setting, and its independence from source data, enabling effective online embedding estimation.
@@ 4.1 H-embedding
Nevertheless, due to its target-centered intrinsicality, the absolute value of H-score is largely dependent on the characteristics of target task, resulting in asymmetry and inconsistent scaling across task pairs. This discrepancy prevents a direct alignment between the reversed H-score and the Euclidean distance among sequentially learned task embeddings. To alleviate this mismatch, we introduce Analytic Hierarchy Process (AHP) normalization (zamir2018taskonomy) to process the H-score in Eqn. 3. Specifically, we construct a pairwise tournament matrix $W^{(j)}\in\mathbb{R}^{j\times j}$ for task $j$, with elements given by:
@@ 4.2.1 Hypernet-Based CL Framework
    Continual learning loss (following von2020continual), to prevent CF by ensuring that given previous task embeddings $\{e^{(n)}\}_{n=1}^{j-1}$, the network weights output by the hypernet before and after the training on task $j$ are analogous
@@ 5.1 Experimental Settings
To comprehensively verify the effectiveness of our framework and further analyze its reliability, we select three representative benchmarks from previous work on CL and perform extensive experiments on them: PermutedMNIST (10 tasks) (goodfellow2013empirical), CIFAR-100 (10 tasks) (krizhevsky2009learning), DomainNet (5 tasks) (peng2019moment) and ImageNet-R (5 tasks, 10 tasks and 20 tasks) (hendrycks2021many) [^6] Training specifics of our experimental studies are listed in Appendix A.6, with codes available at https://github.com/viki760/H-embedding-Guided-Hypernet.. A detailed description of these benchmarks is listed in Appendix A.5.
@@ 5.2 Performance Evaluation
Our primary evaluation study is conducted on the CIFAR-100 and ImageNet-R benchmarks, both consists of 10 tasks. To ensure fairness and comprehensiveness in comparison, a ResNet-32 (he2016deep) and a pretrained ViT-B/16 (dosovitskiy2020image) are selected as the shared backbone models for the full-model and parameter-efficient methods respectively. All reported results are either obtained by our own implementation (averaged over three random seeds with standard deviation) or directly cited from previous works using the same benchmark and backbone.
The choice of baselines is based on the requirements that they should both conform to a rehearsal-free setting and be applicable to these benchmarks and backbones. For a thorough comparison with existing methods to the greatest extent possible, we select representative baselines of varied methodology categories. Full-model strategies include: Regularization Methods: LwF (li2017learning), EWC (kirkpatrick2017overcoming), L2, PredKD+FeatKD (smith2023closer); Architecture Methods: PackNet (mallya2018packnet), HyperNet (von2020continual), WSN (kang2022forget). Parameter-efficient finetuning (PEFT) based strategies include: Prompt Methods: L2P (wang2022learning), DualPrompt (wang2022dualprompt), CODA-Prompt (smith2023coda), HiDe-Prompt (wang2023hierarchical); LoRA Methods: InfLoRA (liang2024inflora), SD-LoRA (wu2025sd). Full finetuning is also included as a basic reference in both backbones. For our framework, we construct full model and LoRA generation versions on ResNet and ViT respectively. The experimental results are summarized in Tab. 1, where our method achieves the highest FAA across both backbones and benchmarks, demonstrating superior overall CL performance. In addition, it attains a competitive DAA that closely aligns with FAA, suggesting strong forward and backward transfer capabilities.
@@ A.1.1 Rehearsal-Free CL
The strictness of CL settings varies with the extent of allowed previous data accessibility. Multi-task learning [caruana1997multitask], with full data availability of all tasks, can actually be viewed as a special case of CL, while rehearsal-free CL [smith2023closer], with no previous data involved in the training of new tasks, is the strictest CL setting under this criteria. Despite the success of rehearsal-based methods in various benchmarks [bang2021rainbow, shin2017continual, belouadah2019il2m], rehearsal-free CL is catching the attention of researchers recently [smith2023closer] because of its low dependency on revisiting previous tasks and therefore broader application in the era of growing data privacy concerns. Existing works on rehearsal-free CL are mostly based on regularization strategies. EWC [kirkpatrick2017overcoming] and SI [zenke2017continual] introduce penalties to restrict the alteration of parameters vital for addressing prior tasks, thereby reducing the risk of CF. LwF [li2017learning, rebuffi2017icarl] proposes a cross-entropy loss between the predicted class distribution of the (n-1)-th task, as generated by the model before and after learning the n-th task. smith2023closer reviews these methods and proposes regularization combinations for better CL performance. In this work, we follow these works and focus on the more challenging rehearsal-free CL setting.

### v4

@@ Abstract
Continual learning (CL) has been a critical topic in contemporary deep neural network applications, where higher levels of both forward and backward transfer are desirable for an effective CL performance. Existing CL strategies primarily focus on task models — either by regularizing model updates or by separating task-specific and shared components — while often overlooking the potential of leveraging inter-task relationships to enhance transfer. To address this gap, we propose a transferability-aware task embedding, termed H-embedding, and construct a hypernet framework under its guidance to learn task-conditioned model weights for CL tasks. Specifically, H-embedding is derived from an information theoretic measure of transferability and is designed to be online and easy to compute. Our method is also characterized by notable practicality, requiring only the storage of a low-dimensional task embedding per task and supporting efficient end-to-end training. Extensive evaluations on benchmarks including CIFAR-100, ImageNet-R, and DomainNet show that our framework performs prominently compared to various baseline and SOTA approaches, demonstrating strong potential in capturing and utilizing intrinsic task relationships. Our code is publicly available at https://github.com/viki760/Hembedding_Guided_Hypernet.
@@ 1 Introduction
Continual learning (CL), also referred to as incremental learning or life-long learning, has become an essential topic in the modern application of deep neural networks. In CL settings, a model is expected to sequentially learn a series of tasks (typically involving either category shifts or data distribution changes (Qu et al., 2021)) to progressively enhance its capabilities (Wang et al., 2024). A key practical challenge in this process is catastrophic forgetting (CF) (Kirkpatrick et al., 2017), a phenomenon where learning a new task undermines the knowledge acquired from previous ones.
@@
The mitigation of CF has long been a primary focus in CL research (Wang et al., 2024) due to its severe hindrance to the overall accumulation of model capacity. Existing approaches can be broadly categorized into three classes: rehearsal-based methods (e.g., ER (Robins, 1995), DER (Buzzega et al., 2020)) maintain a memory buffer of previous samples for later replay; regularization-based methods (e.g., LwF (Li and Hoiem, 2017), SI (Zenke et al., 2017)) constrain parameter updates to protect prior knowledge; and architecture-based methods (e.g., PackNet (Mallya and Lazebnik, 2018), WSN (Kang et al., 2022)) allocate task-specific and task-sharing components of the model from different architecture levels.
@@
However, strategies that explicitly alleviate forgetting often compromise the acquisition of new knowledge, prioritizing retention over adaptability. In contrast, an effective CL system should facilitate the intrinsic knowledge transfer across all tasks, improving future task performance via forward transfer, and preserving (or even enhancing) prior task performance via backward transfer (von Oswald et al., 2020). Moreover, most existing works adopt a model-centric perspective [^1] More discussion in Sec. 2., relying on posterior information obtained during training to manage cross-task interactions. Yet in CL settings, task sequences often exhibit certain degrees of noise and diversity, making such approaches sensitive to sample- and task-level variations. This can lead to potential negative transfer, reduced robustness, and limited scalability, particularly as the number or complexity of tasks increases.
@@
These observations point to a fundamental question in continual learning: Robust and positive knowledge transfers across tasks rely on capturing the underlying task relationships, but how do we efficiently learn and utilize such relations in CL settings? Despite the valuable contributions of previous explorations (e.g. Jin and Kim, 2022; Riemer et al., 2018), they mostly remain constrained by a posterior model-centric paradigm with its associated drawbacks, leaving the prior exploitation of task relations largely underexplored. Based on this insight, we argue that the incorporation of a prior, task-relation-aware guidance can effectively leverage task relation information and help mitigate instability and performance degradation in CL, especially over long or challenging task sequences.
@@
Therefore, in this work we explicitly bring focus to the prior exploitation of task relationships by introducing a statistically grounded, task-relation-aware embedding, and propose a CL framework under its guidance. Specifically, recognizing the shared goal between identifying inter-task relationships and the role of transferability metrics (Ding et al., 2024) in evaluating source-target task compatibility, we propose an online embedding scheme named H-embedding, which distills task-level transferability into a low-dimensional representation through optimization before task training. H-embedding can be learned efficiently without revisiting previous samples by maximizing the consistency between the Euclidean distances among embeddings and the H-score transferability (Bao et al., 2019) among corresponding tasks. To ensure alignment between embedding distances and transferability scores, we apply analytic hierarchy process (AHP) normalization, which significantly improves learning stability over long task sequences while maintaining efficiency. Built upon this embedding, we present a hypernetwork-based CL framework, where a task-conditioned hypernetwork generates task-specific model weights based on H-embedding modulated task embeddings.
@@ 2.1 Continual Learning Strategies
Up to now, there have been many studies dedicated to CL strategies, most of them involving the rehearsal of previous data to alleviate the knowledge degradation caused by CF (e.g. Robins, 1995; Buzzega et al., 2020). However, growing privacy and data safety concerns have made this solution not always feasible, bringing increased attention to the rehearsal-free CL setting Smith et al. (2023b). Most rehearsal-free methods rely on regularization—either in parameter space (e.g. Kirkpatrick et al., 2017; Zenke et al., 2017) or feature space (e.g. Li and Hoiem, 2017; Rebuffi et al., 2017)—by constraining model updates to preserve past knowledge. While effective in CF mitigation, these methods are often based on assuming task similarity and pose limitations to model adaptability. Architecture-based approaches (Wang et al., 2024) offer an alternative by separating task-specific and shared components at various architectural levels (e.g. Mallya and Lazebnik, 2018; Wortsman et al., 2020; Jin and Kim, 2022). With the rise of pretrained models, parameter efficient finetuning (PEFT) components like prompts (e.g. Wang et al., 2022b; Smith et al., 2023a; Wang et al., 2023) and LoRA modules (e.g. Liang and Li, 2024; Wu et al., 2025) are notably gaining popularity in this paradigm. Nevertheless, the allocation of model parts to different tasks usually comes with scaling problems with the growth of task numbers, potentially resulting in insufficient model capacity or excessive model size growth. Task transferability has recently been explored to enhance CL performance (Ermis et al., 2022). Yet, the current method relies on storing both past models and selected samples, which is incompatible with the rehearsal-free setting. Developing strategies that effectively incorporate task transferability while remaining compatible with the online CL paradigm therefore remains an open question.
@@ 2.2 Transferability Metrics
Task transferability (Zamir et al., 2018) investigates the relationships between tasks and provides an effective method to evaluate and select source tasks in transfer learning. It also plays a crucial role in developing strategies for multi-task learning and meta-learning. For ease of use, previous studies have proposed metrics based on task models and data distributions for a quick estimation of transferability (Ding et al., 2024). H-score (Bao et al., 2019; Ibrahim et al., 2022; Wu et al., 2024) uses an information-theoretic framework to evaluate transferability by solving a maximum correlation problem. NCE (Tran et al., 2019) employs conditional entropy to assess transferability and task difficulty. LEEP score (Nguyen et al., 2020; Agostinelli et al., 2022) offers a more generalized metric, defined by measuring the performance of a classifier developed from source model predictions when applied to the target task. LogME (You et al., 2021) assesses target task accuracy using a formulation integrating all possible linear classifiers derived from source model features. OTCE (Tan et al., 2021, 2024) combines optimal transport with conditional entropy to both estimate the domain and task difference between source and target. These metrics are mostly designed with differed assumptions and source accessibility, with their use applicable to different problem settings.
@@ 3.2 Hypernets
Hypernets (Ha et al., 2017), or hypernetworks, are specialized neural networks that generate weights of another neural network, i.e., the target network. Given a target model or component $f$, a hypernetwork $f_{h}$ produces its parameters $\Theta$ conditioned on an auxiliary input. This input can be naturally interpreted as a task embedding, as it reflects the task-specific information needed to generate appropriate parameters. Denoting the task embedding as $e$ and hypernet parameters as $\Theta_{h}$, we have $\Theta=f_{h}(e,\Theta_{h})$. By leveraging task embeddings, hypernets enable flexible parameter modulation across tasks, offering a mechanism for task adaptation without direct weight sharing or parameter interference. Recently, hypernets have gained recognition as a potent tool in deep learning, proving effective across diverse deep learning tasks (Chauhan et al., 2023), including continual learning (von Oswald et al., 2020), causal inference (Chauhan et al., 2024), domain adaptation (Volk et al., 2022), few-shot learning (Sendera et al., 2023), and reinforcement learning (Sarafian et al., 2021).
@@ 3.3 H-score
H-score is firstly introduced by Huang et al. in 2019 as a metric assessing the informativeness of features for a task. Theoretically derived from the maximal correlation interpretation of deep neural networks, its mathematical foundation roots to the information theory work known as maximal correlation analysis, which originates from the works of Hirschfeld, Gebelein and Renyi (Hirschfeld, 1935; Gebelein, 1941; Rényi, 1959) and has been followed and further explored by a broad spectrum of successive work. The H-score of $f$ with regard to the task casting $X$ to $Y$ is defined as:
@@
with input data $X$, label $Y$ and feature extractor $f(X)$. Subsequent research has extended the application of H-score, establishing it as a metric for transferability and validating its efficiency through extensive experimental evaluations (Bao et al., 2019; Ibrahim et al., 2022). This underscores H-score’s potential for transfer learning and its relevance to related challenges. We have chosen to incorporate H-score into our framework due to its robust theoretical reliability, alignment of assumptions with our problem setting, and its independence from source data, enabling effective online embedding estimation.
@@ 4.1 H-embedding
Nevertheless, due to its target-centered intrinsicality, the absolute value of H-score is largely dependent on the characteristics of target task, resulting in asymmetry and inconsistent scaling across task pairs. This discrepancy prevents a direct alignment between the reversed H-score and the Euclidean distance among sequentially learned task embeddings. To alleviate this mismatch, we introduce Analytic Hierarchy Process (AHP) normalization (Zamir et al., 2018) to process the H-score in Eqn. 3. Specifically, we construct a pairwise tournament matrix $W^{(j)}\in\mathbb{R}^{j\times j}$ for task $j$, with elements given by:
@@ 4.2.1 Hypernet-Based CL Framework
    Continual learning loss (following von Oswald et al. (2020)), to prevent CF by ensuring that given previous task embeddings $\{e^{(n)}\}_{n=1}^{j-1}$, the network weights output by the hypernet before and after the training on task $j$ are analogous
@@ 5.1 Experimental Settings
To comprehensively verify the effectiveness of our framework and further analyze its reliability, we select three representative benchmarks from previous work on CL and perform extensive experiments on them: PermutedMNIST (10 tasks) (Goodfellow et al., 2013), CIFAR-100 (10 tasks) (Krizhevsky et al., 2009), DomainNet (5 tasks) (Peng et al., 2019) and ImageNet-R (5 tasks, 10 tasks and 20 tasks) (Hendrycks et al., 2021) [^6] Training specifics of our experimental studies are listed in Appendix A.6, with codes available at https://github.com/viki760/Hembedding_Guided_Hypernet.. A detailed description of these benchmarks is listed in Appendix A.5.
@@ 5.2 Performance Evaluation
Our primary evaluation study is conducted on the CIFAR-100 and ImageNet-R benchmarks, both consists of 10 tasks. To ensure fairness and comprehensiveness in comparison, a ResNet-32 (He et al., 2016) and a pretrained ViT-B/16 (Dosovitskiy et al., 2020) are selected as the shared backbone models for the full-model and parameter-efficient methods respectively. All reported results are either obtained by our own implementation (averaged over three random seeds with standard deviation) or directly cited from previous works using the same benchmark and backbone.
The choice of baselines is based on the requirements that they should both conform to a rehearsal-free setting and be applicable to these benchmarks and backbones. For a thorough comparison with existing methods to the greatest extent possible, we select representative baselines of varied methodology categories. Full-model strategies include: Regularization Methods: LwF (Li and Hoiem, 2017), EWC (Kirkpatrick et al., 2017), L2, PredKD+FeatKD (Smith et al., 2023b); Architecture Methods: PackNet (Mallya and Lazebnik, 2018), HyperNet (von Oswald et al., 2020), WSN (Kang et al., 2022). Parameter-efficient finetuning (PEFT) based strategies include: Prompt Methods: L2P (Wang et al., 2022b), DualPrompt (Wang et al., 2022a), CODA-Prompt (Smith et al., 2023a), HiDe-Prompt (Wang et al., 2023); LoRA Methods: InfLoRA (Liang and Li, 2024), SD-LoRA (Wu et al., 2025). Full finetuning is also included as a basic reference in both backbones. For our framework, we construct full model and LoRA generation versions on ResNet and ViT respectively. The experimental results are summarized in Tab. 1, where our method achieves the highest FAA across both backbones and benchmarks, demonstrating superior overall CL performance. In addition, it attains a competitive DAA that closely aligns with FAA, suggesting strong forward and backward transfer capabilities.
@@ A.1.1 Rehearsal-Free CL
The strictness of CL settings varies with the extent of allowed previous data accessibility. Multi-task learning [Caruana, 1997], with full data availability of all tasks, can actually be viewed as a special case of CL, while rehearsal-free CL [Smith et al., 2023b], with no previous data involved in the training of new tasks, is the strictest CL setting under this criteria. Despite the success of rehearsal-based methods in various benchmarks [Bang et al., 2021, Shin et al., 2017, Belouadah and Popescu, 2019], rehearsal-free CL is catching the attention of researchers recently [Smith et al., 2023b] because of its low dependency on revisiting previous tasks and therefore broader application in the era of growing data privacy concerns. Existing works on rehearsal-free CL are mostly based on regularization strategies. EWC [Kirkpatrick et al., 2017] and SI [Zenke et al., 2017] introduce penalties to restrict the alteration of parameters vital for addressing prior tasks, thereby reducing the risk of CF. LwF [Li and Hoiem, 2017, Rebuffi et al., 2017] proposes a cross-entropy loss between the predicted class distribution of the (n-1)-th task, as generated by the model before and after learning the n-th task. Smith et al. [2023b] reviews these methods and proposes regularization combinations for better CL performance. In this work, we follow these works and focus on the more challenging rehearsal-free CL setting.